# textwidth=0.8textheight=0.9

### v1

@@ 4.3 Comparison with Data-driven MLPs
Figure 5: Performance comparison of cKAN and MLP architectures for data-driven modeling of sand elasto-plasticity.(a) Network architectures (serial and parallel). The models take as input a vector of stress $\bm{\sigma}$, strain $\bm{\varepsilon}$, plastic strain $\bm{\varepsilon}^{\text{p}}$, void ratio $e$, and the incremental strain $\Delta\bm{\varepsilon}\}$, and the output captures the corresponding changes in material response, including changes in void ratio $\Delta e$, incremental stress $\Delta\bm{\sigma}$, and incremental plastic strain $\Delta\bm{\varepsilon}^{\text{p}}\}$.(b) Prediction accuracy under a triaxial loading path defined by $\xi=-1000$, $p_{\text{in}}=375$ kPa, and $e_{\text{in}}=0.64$. The absolute errors in predicting the mean effective stress ($p$) and deviatoric stress ($q$) are assessed by comparing model outputs to reference results obtained from high-accuracy numerical integration.(c) Training convergence trends for serial and parallel architectures. The cKAN sub-networks use relatively compact architectures, with 2 layers of 20 neurons (degree 3) for the void ratio, and 3 layers of 20 neurons (degree 4) for both stress and plastic strain outputs. Moreover, the MLP counterparts employ deeper and wider networks, with 2 layers of 45 neurons for the void ratio and 5 layers of 35 neurons for stress and plastic strain, without polynomial basis functions. All panels are reproduced from  [46]
@@
Figure 6: Evolution of NTK eigenvalues during training of the void ratio subnetwork $e$, using parallel and serial architectures introduced in [46]. (a) cKAN architectures. They exhibit more stable spectral distribution. (b) MLP architectures. They show less structured and more dispersed eigenvalue spectra.
@@
To better understand the advantages of KANs in data-driven learning tasks, it is essential to examine how they perform compared to vanilla MLPs. Such a comparison is significant given the widespread use of MLPs as standard function approximators and the growing interest in KANs as interpretable and efficient alternatives. In this section, we compare the two network types based on three fundamental aspects: accuracy, convergence behavior, and training dynamics analyzed through the Neural Tangent Kernel (NTK) [212, 105] framework. Accuracy evaluates how well each model captures the underlying data distribution, while convergence analysis reflects training efficiency and stability. NTK analysis provides further insight into the models’ training behavior by examining the spectral properties of their kernel matrices. These spectral metrics help explain differences in learning dynamics and generalization capacity. Together, these three criteria provide valuable insights into the practical performance of KANs versus MLPs in data-driven settings. Fig. 5 and Fig. 6 illustrate the comparative results, highlighting the relative strengths and weaknesses of the KAN and MLP models across all three criteria. Fig. 5(a) presents the architectural layouts used in [46], where parallel and serial configurations were applied to both KAN and MLP models to predict the elasto-plastic response of sand under unseen loading paths.
@@ 4.3.1 Accuracy Comparison
Assessing model accuracy is fundamental to understanding how effectively learning frameworks replicate complex behaviors in data-driven tasks. Several studies have compared the accuracy of KAN models across different application domains. In time-series forecasting tasks [168], KANs tracked rapid changes in traffic volume more accurately than MLPs, especially during sudden shifts. In image segmentation, U-KAN [84] achieved a better balance between accuracy and efficiency compared to U-Net. In addition, FC-KAN [192] outperformed MLP models by combining basis functions such as Difference of Gaussians (DoG) and B-splines, with the FC-KAN (DoG+BS) variant achieving the highest accuracy using a quadratic output representation. The study in [46] also examined accuracy by evaluating model predictions during the second stage, where trained networks forecast material responses under unseen loading paths. Fig. 5(b) illustrates the absolute error profiles for key stress components, highlighting differences between cKAN and MLP models in both serial and parallel architectures. Across all configurations, cKAN models consistently demonstrate notably lower errors than their MLP counterparts, indicating superior fidelity in representing nonlinear elasto-plastic behavior. Specifically, serial cKAN architectures reduce prediction errors by a substantial margin compared to serial MLPs, while parallel cKANs also show meaningful improvements over parallel MLPs. These findings suggest that the structured functional form employed by cKANs enhances their ability to accurately model complex stress-strain relationships, supporting their promise for data-driven constitutive modeling in geomechanics and related fields.
@@ 4.3.2 Convergence Analysis
Analyzing convergence patterns offers valuable insight into how different models utilize training iterations to refine their predictions. Several studies have examined convergence behavior in KAN-based models. For example, BSRBF-KAN [205], which combines B-splines and radial basis functions, showed faster convergence compared to MLPs in various learning tasks. Similarly, FC-KAN [192], based on function combinations in KANs, consistently achieved lower training losses than MLPs, attributed to its fast convergence properties. Convergence was also analyzed in [46], where parallel and serial models were trained as part of a two-stage approach. Fig. 5(c) shows the evolution of the cost function during the training stage. The results indicate that cKAN models, in both configurations, exhibit a gradual and sustained decrease in the cost function, reflecting stable and ongoing learning. In contrast, MLP models tend to converge more quickly and often plateau early, particularly in the serial setup, suggesting premature stagnation. While cKANs display some variability near convergence, especially in the parallel case, these fluctuations occur around a consistently declining trend, indicating active adaptation rather than instability. MLPs, however, show abrupt loss reduction followed by flat regions, which implies limited further learning. These convergence dynamics suggest that cKANs are better suited for extended training and offer more reliable long-term improvement, which may ultimately lead to stronger generalization capabilities.
@@ 4.3.3 Spectral Behavior
To compare the spectral behavior of cKAN and MLP, we use the NTK framework to track the evolution of the eigenvalue spectrum during training. NTK theory, introduced by Jacot et al. [212], describes how infinitely wide neural networks evolve under gradient descent, showing that in this limit, the NTK remains constant throughout training. This allows learning to be studied in function space, where convergence rates are directly determined by the NTK’s eigenvalues, directions associated with larger eigenvalues converge more rapidly. Building on this foundation, Seleznova et al. [213, 214] extended NTK analysis to practical, finite-width networks and demonstrated that the NTK can vary significantly during training, particularly in deep networks or under poor initialization. Their results showed that NTK variability increases with network depth, especially in regimes characterized as chaotic or at the edge of chaos. Most NTK-related studies in purely data-driven settings have focused on analyzing MLPs, while limited attention has been given to KANs. Building on the ideas proposed by Mostajeran & Faroughi [46], Fig. 6 presents the evolution of the NTK eigenvalue spectrum for both cKAN and MLP models during the training of networks tasked with predicting the void ratio $e$, across parallel and serial architectural configurations. The cKAN demonstrates a consistently stable and compact spectrum, characterized by low dispersion and smooth progression over time. In contrast, the MLP exhibits irregular spectral dynamics, with final eigenvalues in some cases falling below their initial values, indicating inefficient propagation of learning networks. These findings underscore the superior spectral conditioning and training stability of cKANs, which contribute to their enhanced convergence and generalization in data-driven modeling tasks.
@@ 5 Physics-informed KANs
Physics-informed Kolmogorov–Arnold Networks extend data-driven KANs by embedding domain knowledge, such as partial differential equations, conservation laws, and boundary conditions, directly into the network architecture and training process [66, 53]. This section provides a structured overview of PIKANs, emphasizing their integration of physical laws into learning-based models. The first part details the architectural components and training strategies that enable PIKANs to incorporate governing equations, boundary conditions (e.g., PDEs), and initial constraints into the learning process. The second part highlights a range of applications in solid mechanics, fluid dynamics, and inverse modeling, demonstrating how PIKANs can resolve complex field behaviors and capture sharp gradients or nonlinear material responses with high fidelity. The third part presents a comparative analysis between PIKANs and conventional PINN frameworks, focusing on accuracy, convergence behavior, and spectral characteristics to evaluate their relative performance in solving physics-based problems.
@@ 5.1 Architecture and Training Strategies
Physics-informed KANs integrate physical constraints into their architecture and training process to ensure physically consistent predictions [53]. Unlike purely data-driven approaches, PIKANs leverage the strong form of the governing equations in the loss to enhance accuracy and generalization [66]. As illustrated in Fig. 7, the PIKAN architecture is designed to approximate the solutions of physical systems governed by PDEs/ODEs or other closure models by mapping spatiotemporal input samples to physically meaningful output quantities. The network takes as input a set of spatiotemporal domain samples, denoted by $\bm{\xi}=(\bm{x},t)$, where $\bm{x}\in\Omega\subset\mathbb{R}^{d}$ represents the spatial coordinates and $t\in[0,T]$ denotes time. The corresponding output represents physically relevant quantities, such as pressure, velocity, displacement, or concentration, depending on the specific application. Let the behavior of the system be governed by a PDE of the general form,
@@
where $u$ is the solution to be approximated by the network, $\mathcal{N}$ is a (possibly nonlinear) differential operator that encodes the governing physical laws, $f$ is a source or forcing term, $u_{0}$ defines the initial state of the system, and $\mathcal{B}$ represents a boundary operator with associated boundary values $g$. To train PIKANs, a composite loss function is used to enforce consistency with observed data and compliance with physical constraints. This loss function is defined as [2, 66],
@@
These constraints may be imposed as either soft or hard. The key difference between soft and hard constraints involves the integration of physical conditions within the learning framework. For soft constraints [66], physical laws and conditions are enforced indirectly by adding penalty terms to the loss function akin to the approach outlined in Eq. (16). This approach encourages the network to approximate the desired behavior but does not guarantee strict satisfaction of the constraints. In contrast, hard constraints are embedded directly into the network architecture [215, 216], ensuring that the output inherently satisfies the initial and boundary conditions. This is typically achieved by re-parameterizing the network output as,
@@
Training PIKANs also involves optimizing the complex, multi-term loss function (Eq. (16)), which requires careful selection of optimization strategies. Although first-order methods such as Adam are widely used, recent research has shown that second-order optimizers, particularly self-scaled Broyden (SSBroyden) [201] and BFGS with Wolfe line search [199, 200, 204], can significantly accelerate convergence and improve accuracy in training PIKANs [107]. These quasi-Newton methods exploit curvature information from the loss landscape, helping to avoid local minima and saddle points that often trap first-order optimizers. A promising approach is to use hybrid optimization strategies, where the network is first trained with adaptive first-order methods like RAdam to quickly reduce the loss, followed by second-order fine-tuning with BFGS variants. This combination has been shown to achieve high accuracy and stable convergence across various problems [102]. The effectiveness of these optimizers also depends on architectural and numerical factors. For example, the use of Chebyshev polynomial-based activation in tanh-cPIKANs [102] enhances optimization by producing smoother gradients, while double-precision arithmetic provides better numerical stability and optimization performance. Additionally, adaptively adjusting the weights of different loss components helps to balance learning signals and avoid domination by any single term during training [217, 218, 219, 220].
@@
PIKAN Used different variants of KAN (B-spline, Chebyshev, Legendre, Hermit, Jacoiy, Bayesian-PIKAN); Applied to Helmholtz, Navier–Stokes, Allen–Cahn, and Reaction–diffusion equations. [66]
@@
Scaled-cPIKAN Combined Chebyshev-based KANs with a scaling approach for spatial variables and governing equations; Applied to PDEs with oscillatory dynamics over large domains, such as diffusion, Helmholtz, Allen–Cahn, and reaction–diffusion equations. [68]
@@
KAN-ODE Proposed an interpretable and modular framework without prior knowledge by integrating KANs into a Neural ODE structure; Used Gaussian RBFs basis functions; Applied to Lotka–Volterra, Fisher-KPP, Burgers’, Schrödinger, and Allen-Cahn equations. [221]
@@
HRKANs Enhanced the approximation capability of KANs by introducing higher-order ReLU basis functions in place of linear splines; Applied to the linear Poisson equation and the nonlinear Burgers’ equation with viscosity. [222]
@@
SPIKAN Introduced a separable architecture for PIKANs by applying the principle of variable separation; Used B-splines as basis functions; Applied to 2D Helmholtz, 2D steady lid-driven cavity flow, 1D+1 Allen–Cahn, and 2D+1 Klein–Gordon equations [101]
@@
KKAN Designed a two-block architecture with MLP-based inner functions and basis-function-based outer functions to form a universal approximator; Used RBFs as outer blocks; Applied to the Allen–Cahn equation. [117]
@@
Bayesian PIKANs Proposed a gradient-free method combining DTEKI with Chebyshev KANs, with incorporation of active subspaces method; Applied to transport, 1D diffusion, 1D nonlinear and 2D Darcy flow equations. [223]
@@
LeanKAN Proposed a direct and modular alternative to MultKAN and AddKAN layers to reduce parameters and hyperparameters; Used RBF expansions combined with a base activation function; Applied to the classical Lotka–Volterra predator–prey model. [224]
@@
asKAN Proposed active subspace embedded KAN to circumvent the inflexibility and organized KANs in a hierarchical framework; Used a third-order spline function, with 5 grid points; Applied to the Poisson equation and scattering sound field reconstruction. [225]
@@
Research on architectural advancements in PIKANs has led to various structural modifications tailored to specific tasks. A non-exhaustive overview of these developments is provided in Table 3. A key line of work is presented in [66], where extensive evaluations of different basis functions were conducted. In this study, the B-spline activations, commonly used in early KAN architectures, were replaced with orthogonal polynomial functions such as Chebyshev, Jacobi, and Legendre. These replacements were further improved through the incorporation of residual-based attention and entropy viscosity methods. To improve scalability, new variants such as DF-PIKANs and DD-PIKANs [103] were introduced. These models leverage wavelet and B-spline activations to reduce model complexity while preserving accuracy across a wide range of PDEs, particularly those dominated by convection. Building on this, SPIKAN [101] employed a separation-of-variables strategy, effectively reducing both training cost and memory usage in high-dimensional scenarios. Furthermore, HRKANs [222] adopted Higher-order-ReLU activations, aiming to streamline implementation and enhance solver performance for problems such as Poisson and Burgers’ equations.

### v2

@@ 4.3 Comparison with Data-driven MLPs
Figure 5: Performance comparison of cKAN and MLP architectures for data-driven modeling of sand elasto-plasticity.(a) Network architectures (serial and parallel). The models take as input a vector of stress $\bm{\sigma}$, strain $\bm{\varepsilon}$, plastic strain $\bm{\varepsilon}^{\text{p}}$, void ratio $e$, and the incremental strain $\Delta\bm{\varepsilon}\}$, and the output captures the corresponding changes in material response, including changes in void ratio $\Delta e$, incremental stress $\Delta\bm{\sigma}$, and incremental plastic strain $\Delta\bm{\varepsilon}^{\text{p}}\}$.(b) Prediction accuracy under a triaxial loading path defined by $\xi=-1000$, $p_{\text{in}}=375$ kPa, and $e_{\text{in}}=0.64$. The absolute errors in predicting the mean effective stress ($p$) and deviatoric stress ($q$) are assessed by comparing model outputs to reference results obtained from high-accuracy numerical integration.(c) Training convergence trends for serial and parallel architectures. The cKAN sub-networks use relatively compact architectures, with 2 layers of 20 neurons (degree 3) for the void ratio, and 3 layers of 20 neurons (degree 4) for both stress and plastic strain outputs. Moreover, the MLP counterparts employ deeper and wider networks, with 2 layers of 45 neurons for the void ratio and 5 layers of 35 neurons for stress and plastic strain, without polynomial basis functions. All panels present new analysis based on mostajeran2024epi .
@@
Figure 6: Evolution of NTK eigenvalues during training of the void ratio subnetwork $e$, using parallel and serial architectures introduced in mostajeran2024epi . (a) cKAN architectures. They exhibit more stable spectral distribution. (b) MLP architectures. They show less structured and more dispersed eigenvalue spectra. All panels present new analysis based on mostajeran2024epi .
@@
To better understand the advantages of KANs in data-driven learning tasks, it is essential to examine how they perform compared to vanilla MLPs. Such a comparison is significant given the widespread use of MLPs as standard function approximators and the growing interest in KANs as interpretable and efficient alternatives. In this section, we compare the two network types based on three fundamental aspects: accuracy, convergence behavior, and training dynamics analyzed through the Neural Tangent Kernel (NTK) jacot2018neural , faroughi2025neural framework. Accuracy evaluates how well each model captures the underlying data distribution, while convergence analysis reflects training efficiency and stability. NTK analysis provides further insight into the models’ training behavior by examining the spectral properties of their kernel matrices. These spectral metrics help explain differences in learning dynamics and generalization capacity. Together, these three criteria provide valuable insights into the practical performance of KANs versus MLPs in data-driven settings. Fig. 5 and Fig. 6 illustrate the comparative results, highlighting the relative strengths and weaknesses of the KAN and MLP models across all three criteria. Fig. 5(a) presents the architectural layouts used in mostajeran2024epi , where parallel and serial configurations were applied to both KAN and MLP models to predict the elasto-plastic response of sand under unseen loading paths.
@@ 4.3.1 Accuracy Comparison
Assessing model accuracy is fundamental to understanding how effectively learning frameworks replicate complex behaviors in data-driven tasks. Several studies have compared the accuracy of KAN models across different application domains. In comparative experiments, the P1-KAN warin2024p1 achieved higher accuracy than multilayer perceptrons and other KAN variants for both smooth and highly irregular functions, demonstrating strong robustness across different regularity regimes. In time-series forecasting tasks vaca2024kolmogorov , KANs tracked rapid changes in traffic volume more accurately than MLPs, especially during sudden shifts. In image segmentation, U-KAN li2025u achieved a better balance between accuracy and efficiency compared to U-Net. In addition, FC-KAN ta2024fc outperformed MLP models by combining basis functions such as Difference of Gaussians (DoG) and B-splines, with the FC-KAN (DoG+BS) variant achieving the highest accuracy using a quadratic output representation. The study in mostajeran2024epi also examined accuracy by evaluating model predictions during the second stage, where trained networks forecast material responses under unseen loading paths. Fig. 5(b) illustrates the absolute error profiles for key stress components, highlighting differences between cKAN and MLP models in both serial and parallel architectures. Across all configurations, cKAN models consistently demonstrate notably lower errors than their MLP counterparts, indicating superior fidelity in representing nonlinear elasto-plastic behavior. Specifically, serial cKAN architectures reduce prediction errors by a substantial margin compared to serial MLPs, while parallel cKANs also show meaningful improvements over parallel MLPs. These findings suggest that the structured functional form employed by cKANs enhances their ability to accurately model complex stress-strain relationships, supporting their promise for data-driven constitutive modeling in geomechanics and related fields.
@@ 4.3.2 Convergence Analysis
Analyzing convergence patterns offers valuable insight into how different models utilize training iterations to refine their predictions. Several studies have examined convergence behavior in KAN-based models. For example, BSRBF-KAN ta2024bsrbf , which combines B-splines and radial basis functions, showed faster convergence compared to MLPs in various learning tasks. Similarly, FC-KAN ta2024fc , based on function combinations in KANs, consistently achieved lower training losses than MLPs, attributed to its fast convergence properties. The P1-KAN warin2024p1 also exhibited stable and consistent convergence behavior, with smooth error reduction during training, particularly in stochastic optimization problems such as hydraulic valley modeling. Convergence was also analyzed in mostajeran2024epi , where parallel and serial models were trained as part of a two-stage approach. Fig. 5(c) shows the evolution of the cost function during the training stage. The results indicate that cKAN models, in both configurations, exhibit a gradual and sustained decrease in the cost function, reflecting stable and ongoing learning. In contrast, MLP models tend to converge more quickly and often plateau early, particularly in the serial setup, suggesting premature stagnation. While cKANs display some variability near convergence, especially in the parallel case, these fluctuations occur around a consistently declining trend, indicating active adaptation rather than instability. MLPs, however, show abrupt loss reduction followed by flat regions, which implies limited further learning. These convergence dynamics suggest that cKANs are better suited for extended training and offer more reliable long-term improvement, which may ultimately lead to stronger generalization capabilities.
@@ 4.3.3 Spectral Behavior
To compare the spectral behavior of cKAN and MLP, we use the NTK framework to track the evolution of the eigenvalue spectrum during training. NTK theory, introduced by Jacot et al. jacot2018neural , describes how infinitely wide neural networks evolve under gradient descent, showing that in this limit, the NTK remains constant throughout training. This allows learning to be studied in function space, where convergence rates are directly determined by the NTK’s eigenvalues, directions associated with larger eigenvalues converge more rapidly. Building on this foundation, Seleznova et al. seleznova2022analyzing , seleznova2022neural extended NTK analysis to practical, finite-width networks and demonstrated that the NTK can vary significantly during training, particularly in deep networks or under poor initialization. Their results showed that NTK variability increases with network depth, especially in regimes characterized as chaotic or at the edge of chaos. Most NTK-related studies in purely data-driven settings have focused on analyzing MLPs, while limited attention has been given to KANs. Building on the ideas proposed by Mostajeran & Faroughi mostajeran2024epi , Fig. 6 presents the evolution of the NTK eigenvalue spectrum for both cKAN and MLP models during the training of networks tasked with predicting the void ratio $e$, across parallel and serial architectural configurations. In all experiments, the networks operate in the finite-width regime, where each layer contains a limited number of neurons and the weights evolve throughout training. As a result, the NTK is not constant but evolves dynamically over time, capturing feature learning effects rather than purely kernel regression behavior. The cKAN demonstrates a consistently stable and compact spectrum, characterized by low dispersion and smooth progression over time. In contrast, the MLP exhibits irregular spectral dynamics, with final eigenvalues in some cases falling below their initial values, indicating inefficient propagation of learning networks. Rigas et al. rigas2025initialization also emphasized the link between initialization and NTK dynamics in spline-based KANs. They showed that baseline schemes often produce spectra that collapse during training, leading to poor conditioning, whereas Glorot- and power-law-based strategies yield eigenvalue spectra that stabilize early and remain well-conditioned. These insights reinforce the importance of initialization in maintaining stable NTK behavior, in line with the spectral stability observed for KANs. These findings underscore the superior spectral conditioning and training stability of cKANs, which contribute to their enhanced convergence and generalization in data-driven modeling tasks.
@@ 5 Physics-informed KANs
Physics-informed Kolmogorov–Arnold Networks extend data-driven KANs by embedding domain knowledge, such as partial differential equations, conservation laws, and boundary conditions, directly into the network architecture and training process shukla2024comprehensive , wang2024kolmogorov . This section provides a structured overview of PIKANs, emphasizing their integration of physical laws into learning-based models. The first part details the architectural components and training strategies that enable PIKANs to incorporate governing equations, boundary conditions (e.g., PDEs), and initial constraints into the learning process. The second part highlights a range of applications in solid mechanics, fluid dynamics, and inverse modeling, demonstrating how PIKANs can resolve complex field behaviors and capture sharp gradients or nonlinear material responses with high fidelity. The third part presents a comparative analysis between PIKANs and conventional PINN frameworks, focusing on accuracy, convergence behavior, and spectral characteristics to evaluate their relative performance in solving physics-based problems.
@@ 5.1 Architecture and Training Strategies
Physics-informed KANs integrate physical constraints into their architecture and training process to ensure physically consistent predictions wang2024kolmogorov . Unlike purely data-driven approaches, PIKANs leverage the strong form of the governing equations in the loss to enhance accuracy and generalization shukla2024comprehensive . As illustrated in Fig. 7, the PIKAN architecture is designed to approximate the solutions of physical systems governed by PDEs/ODEs or other closure models by mapping spatiotemporal input samples to physically meaningful output quantities. The network takes as input a set of spatiotemporal domain samples, denoted by $\bm{\xi}=(\bm{x},t)$, where $\bm{x}\in\Omega\subset\mathbb{R}^{d}$ represents the spatial coordinates and $t\in[0,T]$ denotes time. The corresponding output represents physically relevant quantities, such as pressure, velocity, displacement, or concentration, depending on the specific application. Let the behavior of the system be governed by a PDE of the general form,
@@
where $u$ is the solution to be approximated by the network, $\mathcal{N}$ is a (possibly nonlinear) differential operator that encodes the governing physical laws, $f$ is a source or forcing term, $u_{0}$ defines the initial state of the system, and $\mathcal{B}$ represents a boundary operator with associated boundary values $g$. To train PIKANs, a composite loss function is used to enforce consistency with observed data and compliance with physical constraints. This loss function is defined as raissi2019physics , shukla2024comprehensive ,
@@
These constraints may be imposed as either soft or hard. The key difference between soft and hard constraints involves the integration of physical conditions within the learning framework. For soft constraints shukla2024comprehensive , physical laws and conditions are enforced indirectly by adding penalty terms to the loss function akin to the approach outlined in Eq. (16). This approach encourages the network to approximate the desired behavior but does not guarantee strict satisfaction of the constraints. In contrast, hard constraints are embedded directly into the network architecture lu2021physics , lai2025hard , ensuring that the output inherently satisfies the initial and boundary conditions. This is typically achieved by re-parameterizing the network output as,
@@
Training PIKANs also involves optimizing the complex, multi-term loss function (Eq. (16)), which requires careful selection of optimization strategies. Although first-order methods such as Adam are widely used, recent research has shown that second-order optimizers, particularly self-scaled Broyden (SSBroyden) al2014broyden and BFGS with Wolfe line search nawi2006improved , saputro2017limited , liu1989limited , can significantly accelerate convergence and improve accuracy in training PIKANs kiyani2025optimizer . These quasi-Newton methods exploit curvature information from the loss landscape, helping to avoid local minima and saddle points that often trap first-order optimizers. A promising approach is to use hybrid optimization strategies, where the network is first trained with adaptive first-order methods like RAdam to quickly reduce the loss, followed by second-order fine-tuning with BFGS variants. This combination has been shown to achieve high accuracy and stable convergence across various problems daryakenari2025representation . The effectiveness of these optimizers also depends on architectural and numerical factors. For example, the use of Chebyshev polynomial-based activation in tanh-cPIKANs daryakenari2025representation enhances optimization by producing smoother gradients, while double-precision arithmetic provides better numerical stability and optimization performance. Additionally, adaptively adjusting the weights of different loss components helps to balance learning signals and avoid domination by any single term during training liu2022physics , gao2025physics , heydari2019softadapt , barron2019general .
@@
PIKAN Used different variants of KAN (B-spline, Chebyshev, Legendre, Hermit, Jacoiy, Bayesian-PIKAN); Applied to Helmholtz, Navier–Stokes, Allen–Cahn, and Reaction–diffusion equations. shukla2024comprehensive
@@
Scaled-cPIKAN Combined Chebyshev-based KANs with a scaling approach for spatial variables and governing equations; Applied to PDEs with oscillatory dynamics over large domains, such as diffusion, Helmholtz, Allen–Cahn, and reaction–diffusion equations. mostajeran2025scaled
@@
KAN-ODE Proposed an interpretable and modular framework without prior knowledge by integrating KANs into a Neural ODE structure; Used Gaussian RBFs basis functions; Applied to Lotka–Volterra, Fisher-KPP, Burgers’, Schrödinger, and Allen-Cahn equations. koenig2024kan
@@
HRKANs Enhanced the approximation capability of KANs by introducing higher-order ReLU basis functions in place of linear splines; Applied to the linear Poisson equation and the nonlinear Burgers’ equation with viscosity. so2024higher
@@
SPIKAN Introduced a separable architecture for PIKANs by applying the principle of variable separation; Used B-splines as basis functions; Applied to 2D Helmholtz, 2D steady lid-driven cavity flow, 1D+1 Allen–Cahn, and 2D+1 Klein–Gordon equations jacob2024spikans
@@
KKAN Designed a two-block architecture with MLP-based inner functions and basis-function-based outer functions to form a universal approximator; Used RBFs as outer blocks; Applied to the Allen–Cahn equation. toscano2025kkans
@@
Bayesian PIKANs Proposed a gradient-free method combining DTEKI with Chebyshev KANs, with incorporation of active subspaces method; Applied to transport, 1D diffusion, 1D nonlinear and 2D Darcy flow equations. gao2025scalable
@@
LeanKAN Proposed a direct and modular alternative to MultKAN and AddKAN layers to reduce parameters and hyperparameters; Used RBF expansions combined with a base activation function; Applied to the classical Lotka–Volterra predator–prey model. koenig2025
@@
asKAN Proposed active subspace embedded KAN to circumvent the inflexibility and organized KANs in a hierarchical framework; Used a third-order spline function, with 5 grid points; Applied to the Poisson equation and scattering sound field reconstruction. zhou2025
@@
Research on architectural advancements in PIKANs has led to various structural modifications tailored to specific tasks. A non-exhaustive overview of these developments is provided in Table 3. A key line of work is presented in shukla2024comprehensive , where extensive evaluations of different basis functions were conducted. In this study, the B-spline activations, commonly used in early KAN architectures, were replaced with orthogonal polynomial functions such as Chebyshev, Jacobi, and Legendre. These replacements were further improved through the incorporation of residual-based attention and entropy viscosity methods. To improve scalability, new variants such as DF-PIKANs and DD-PIKANs patra2024physics were introduced. These models leverage wavelet and B-spline activations to reduce model complexity while preserving accuracy across a wide range of PDEs, particularly those dominated by convection. Building on this, SPIKAN jacob2024spikans employed a separation-of-variables strategy, effectively reducing both training cost and memory usage in high-dimensional scenarios. Furthermore, HRKANs so2024higher adopted Higher-order-ReLU activations, aiming to streamline implementation and enhance solver performance for problems such as Poisson and Burgers’ equations.